# A new class of positive linear operators preserving logarithmic functions

### v1

@@ 5 Further approximation properties
In this section, we explore the approximation properties of the operators $(\mathscr{L}_{n})_{n}$, using their connection with the classical King-type operators ([27, 19]). In particular, we derive a quantitative estimate for the approximation error and we conclude with a Voronovskaja-type asymptotic result.
@@
$\mathcal{V}_{n}(e_{0},x)=1,\quad e_{0}(x)\equiv 1,$ (11)
$\mathcal{V}_{n}(e_{1},x)=r_{n}(x),\quad e_{1}(x)\equiv x,$ (12)
$\mathcal{V}_{n}(e_{2},x)=\frac{r_{n}(x)}{n}+\frac{n-1}{n}r_{n}^{2}(x),\quad e_{2}(x)\equiv x^{2}.$ (13)
Furthermore, he showed that $\mathcal{V}_{n}(f,x)\to f(x)$ as $n\to+\infty$, for every $f\in C([0,1])$ and $x\in[0,1]$, if and only if $\lim_{n\to+\infty}r_{n}(x)=x$.
@@
By means of this relation, it is possible to provide a quantitative estimate of the error of approximation for the operators $(\mathscr{L}_{n})_{n}$. In this regard, we state the following theorem.
@@
for $n\in\mathbb{N}$, where $\gamma_{n}:=\max_{x\in[0,1]}(a_{n}(x)-x).$
@@
$\leq\left\{\frac{\|a_{n}\|_{\infty}\max_{x\in[0,1]}|1-a_{n}(x)|}{n\delta^{2}}+\frac{1}{\delta^{2}}\max_{x\in[0,1]}(a_{n}(x)-x)^{2}\right\}^{\frac{1}{2}}.$
@@
Denoting $\gamma_{n}:=\max_{x\in[0,1]}(a_{n}(x)-x)$, so that
@@
In order to ensure that the last term of the estimate (17) does not diverge as $n\to+\infty$, we analyze the asymptotic behavior of the quantity $\sqrt{n}\gamma_{n}$, where $\gamma_{n}=\displaystyle\max_{x\in[0,1]}(a_{n}(x)-x)$. To this end, we recall the notation (6), i.e.,
@@
so that $\varepsilon_{n}\to 0^{+}$ as $n\to+\infty$. Accordingly, we rewrite
@@
Therefore,
$a_{n}(x)-x=\varepsilon_{n}\left(\frac{x-x^{2}}{2}\right)+o(\varepsilon_{n}),$
from which
@@
Now, we establish a Voronovskaja-type asymptotic formula for the operators $(\mathscr{L}_{n})_{n}$.
@@
Since the function $\psi(y)$ is bounded and tends to zero as $y\to 0^{+}$, fixed $\varepsilon>0$, there exists $\widetilde{\delta}>0$ such that, for every $y$ with $0<y\leq\widetilde{\delta}$, then $|\psi(y)|<\varepsilon$. As a consequence, we can write
$n|R_{n}(x)|$ $\leq n\Bigg\{\sum_{\left|\frac{k}{n}-a_{n}(x)\right|\leq\frac{\widetilde{\delta}}{2}}+\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\widetilde{\delta}}{2}}\Bigg\}\left(\frac{k}{n}-x\right)^{2}\left|\psi\left(\frac{k}{n}-x\right)\right|\,$
@@
$\left|\frac{k}{n}-x\right|\leq\left|\frac{k}{n}-a_{n}(x)\right|+\left|a_{n}(x)-x\right|\leq\widetilde{\delta}$,
@@
$R_{2}$ $\leq\|\psi\|_{\infty}\cdot n\ \sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\widetilde{\delta}}{2}}\left(\frac{k}{n}-x\right)^{2}\,p_{n,k}(a_{n}(x))$
$\leq\|\psi\|_{\infty}\cdot n\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\widetilde{\delta}}{2}}\left[2\left(\frac{k}{n}-a_{n}(x)\right)^{2}+2\left(a_{n}(x)-x\right)^{2}\right]\,p_{n,k}(a_{n}(x))$
$\leq{2}\|\psi\|_{\infty}\left\{n\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\widetilde{\delta}}{2}}\left(\frac{k}{n}-a_{n}(x)\right)^{2}\,p_{n,k}(a_{n}(x))+n(a_{n}(x)-x)^{2}\right\},$
@@
$\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\widetilde{\delta}}{2}}p_{n,k}(a_{n}(x))\leq C\,n^{-2}.$
@@
${2}\|\psi\|_{\infty}\,n\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\widetilde{\delta}}{2}}\left(\frac{k}{n}-a_{n}(x)\right)^{2}p_{n,k}(a_{n}(x))$
$\leq{2}\|\psi\|_{\infty}\,n\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\widetilde{\delta}}{2}}p_{n,k}(a_{n}(x))\leq{2}\|\psi\|_{\infty}C\,n^{-1}.$
@@ 6 Saturation results and inverse theorems
where $f:[0,1]\to\mathbb{R}$ and $f_{\mu}:=f/\ln_{\mu}$ as usual. By simple calculations, it is easy to see that the above differential operator can be rewritten in the following standard form:
@@
In [18], Garrancho and Cárdenas-Morales, inspired by the so-called parabola technique introduced by Bajsanski and Bojanic in [11] (see also Chapter 5 of [16] for a more detailed explanation), established the following useful result involving differential operators of the form (25).
@@ 7 Shape-preserving properties
In this section we study some shape preserving properties of the linear and positive operators $(\mathscr{L}_{n})_{n}$.
@@
$\displaystyle\left(\frac{\mathscr{L}_{n}f}{\ln_{\mu}}\right)^{\prime\prime}(x)=$ $\hskip 2.84526pta_{n}^{\prime\prime}(x)\sum_{k=0}^{n-1}\Delta_{1}f_{\mu}\left(\frac{k}{n}\right)\,\,p_{n-1,k}(a_{n}(x))$
$+\frac{n-1}{n}a_{n}^{\prime}(x)^{2}\hskip 2.84526pt\sum_{k=0}^{n-2}\Delta_{2}f_{\mu}\left(\frac{k}{n}\right)\,\,p_{n-2,k}(a_{n}(x)).$ (30)
Since $a_{n}(x)$ is increasing on the interval $[0,1]$, it follows from expressions (28) and (29) that, if $f_{\mu}$ is increasing on $[0,1]$, both ${\mathscr{L}_{n}f}/{\ln_{\mu}}$ and $\mathscr{L}_{n}(f,x)$ are increasing on $[0,1]$. And if in addition $f_{\mu}$ is concave then, taking into account that $a_{n}(x)$ is concave, ${\mathscr{L}_{n}f}/{\ln_{\mu}}$ is concave as well.
@@
and $\mathcal{P}$ denotes the family of all the partitions of the interval $[0,1]$, that is, sets of the form $P=\{0=x_{0}<x_{1}<\cdots<x_{m}=1\}$. Furthermore, we recall that $BV_{\mu}([0,1])$ can be endowed with the norm
@@
Now, since the expressions (2) and (28) hold, we obtain that
@@
(a) If $f_{\mu}$ is increasing and convex, then, for every $n\in\mathbb{N}$ and $x\in]0,1[$,
@@
(b) If $f_{\mu}$ is decreasing and concave, then, for every $n\in\mathbb{N}$ and $x\in]0,1[$,
@@
We will prove part $(a)$ since part $(b)$ follows immediately applying the same considerations of $(a)$ to the function f-f, so that $f_{\mu}$ will be increasing and convex.
@@
$\mathscr{L}_{n}(f,x)=\ln_{\mu}(x)\cdot B_{n}\left(f_{\mu},a_{n}(x)\right),$
where $f_{\mu}:=f/\ln_{\mu}$. Since $f_{\mu}$ is convex on $[0,1]$, as a consequence of the classical shape preserving properties of Bernstein operators (see [17]), we have that
@@
for every $y\in[0,1]$. Moreover, since $a_{n}(x)\geq a_{n+1}(x)$ for every $x\in[0,1]$, $n\in\mathbb{N}$, and $f_{\mu}$ is increasing, it follows that
@@
Finally, $\mathscr{L}_{n}(f,x)\to f(x)$, as $n\to+\infty$, on every continuity point $x$ of $f$ (Theorem 4.1), therefore on every $x\in]0,1[$: indeed, $f_{\mu}$ is continuous in ]0,1[]0,1[, since it is convex, and so is $f=f_{\mu}\cdot\ln_{\mu}$. Therefore, since $\mathscr{L}_{n}(f,x)$ is decreasing with respect to $n$, we deduce that
@@ 8 Final remarks and future developments: a simple application to signal denoising
In this paper, we introduced a new definition of positive linear operators, which is characterized by the preservation of the function $\ln_{\mu}$. For such operators, we established a number of approximation properties, such as uniform convergence, Voronovskaja-type formula, saturation and inverse theorems based on the solution of a suitable differential problem, and shape preserving properties.
@@
In real world situations, it can occur that a given signal $f$ is characterized by the presence of multiplicative noise, such as Gaussian multiplicative noise, uniform multiplicative noise, and several others. For instance, a typical example of multiplicative noise is the so-called speckle, that typically affects remote sensing data, or tomography and ultrasound ecography images. Other cases in which multiplicative noise arises is in stochastic models for finance or in wireless signal applications.
@@
Let $f:[0,1]\to\mathbb{R}^{+}$ be a given signal modelling a physical quantity. Suppose that $f$ is sampled (by a certain measurement tool) at a certain time $t>0$ in the following uniform spaced nodes $k/n$, k=0,1,,nk=0,1,...,n, $n\in\mathbb{N}$, obtaining the following sequence of measured sample values $(y_{k,n})_{k}$.
We now suppose that $f$ is affected by a multiplicative noise function of the form: $n(x,t)=1+\mu(t)+x$, where the terms $\mu(t)$ can be viewed as a Gaussian random variable of mean 0 and variance $\sigma^{2}$, namely
@@
Hence, we expect that the sampled values $(y_{k,n})_{k}$ are of the form:
$y_{k,n}\approx\left(1+\mu(t)+\frac{k}{n}\right)\,f(k/n),$
where $t>0$ represents the time in which the samples have been measured.
Let now define $g(x):=\left(1+\mu(t)+x\right)f(x)$, for a fixed $t>0$, and assume that $\mu(t)>0$; thus:
$\ln g\ =\ \ln_{\mu(t)}\,+\,\ln f,$
from which we get:
@@
where
${\mathscr{L}}_{n}(\ln g,x)\ =\ \ln_{\mu}(x)\sum_{k=0}^{n}\frac{\ln y_{k,n}}{\ln_{\mu}\left(\frac{k}{n}\right)}\hskip 2.84526ptp_{n,k}(a_{n}(x)).$
We now present a numerical example. Let us consider, for instance, three values of $\mu$ randomly generated as a Gaussian random variable ${\cal N}(0,\sigma^{2})$, with $\sigma=0.5$:
$\mu_{1}=0.2688,\quad\mu_{2}=0.9169,\quad\mu_{3}=1.1294.$
and let $f:[0,1]\to\mathbb{R}^{+}$ defined as:
@@
We now consider the following noise-free reconstructions of $f$, for $n=10,30$ in Fig.s 2, 3 and 4. In the captions of the figures below the maximum error of approximation is reported.
Figure 2: The (blue) solid line represents the original signal $f$, the dotted (red) plot represents the signal $g$ affected by the multiplicative noise, the asterisk (yellow) plot represent the denoised signal by (31) for $\mu_{1}$, with $n=10$ (on the left) and $n=30$ (on the right). The maximum reconstruction errors are $0.1109$ and $0.0343$, respectively.
Figure 3: The (blue) solid line represents the original signal $f$, the dotted (red) plot represents the signal $g$ affected by the multiplicative noise, the asterisk (yellow) plot represent the denoised signal by (31) for $\mu_{2}$, with $n=10$ (on the left) and $n=30$ (on the right). The maximum reconstruction error are $0.0658$ and $0.0202$, respectively.
Figure 4: The (blue) solid line represents the original signal $f$, the dotted (red) plot represents the signal $g$ affected by the multiplicative noise, the asterisk (yellow) plot represent the denoised signal by (31) for $\mu_{3}$, with $n=10$ (on the left) and $n=30$ (on the right). The maximum reconstruction errors are $0.0622$ and $0.0191$, respectively.
Obviously, the above experiment represents only a toy model which serves to introduce a possible application of the new logarithmic-type operators; the main idea was to exploit the logarithm-preservation property besides the approximation properties of $({\mathscr{L}}_{n})_{n}$ to linearize a nonlinear source of noise. In future works, we aim to improve the above simple model, for instance, in order to deal with multivariate signals to implement a despeckle algorithm for remote sensing images.

### v2

@@ 4 Further approximation properties
In this section, we explore the approximation properties of the operators $\mathscr{L}_{n}$, using their connection with the classical King-type operators ([30, 22, 20]). In particular, we derive a quantitative estimate for the approximation error and we conclude with a Voronovskaja-type asymptotic result.
@@
$\mathcal{V}_{n}(e_{0},x)=1,$ (10)
$\mathcal{V}_{n}(e_{1},x)=r_{n}(x),$ (11)
$\mathcal{V}_{n}(e_{2},x)=\frac{r_{n}(x)}{n}+\frac{n-1}{n}r_{n}^{2}(x).$ (12)
Furthermore, he showed that $\mathcal{V}_{n}(f,x)\to f(x)$ as $n\to+\infty$, for every $f\in C([0,1])$ and $x\in[0,1]$, if and only if $\lim\limits_{n\to+\infty}r_{n}(x)=x$.
@@
By means of this relation, it is possible to provide a quantitative estimate of the error of approximation for the operators $\mathscr{L}_{n}$. In this regard, we state the following theorem.
@@
for $n\in\mathbb{N}$, where $\gamma_{n}:=\max\limits_{x\in[0,1]}(a_{n}(x)-x).$
@@
$\leq\left\{\frac{\|a_{n}\|_{\infty}\max\limits_{x\in[0,1]}|1-a_{n}(x)|}{n\delta^{2}}+\frac{1}{\delta^{2}}\max\limits_{x\in[0,1]}(a_{n}(x)-x)^{2}\right\}^{\frac{1}{2}}.$
@@
Denoting $\gamma_{n}:=\max\limits_{x\in[0,1]}(a_{n}(x)-x)$, so that
@@
For the sake of completeness, we analyze the asymptotic behavior of the quantity $\sqrt{n}\gamma_{n}$, where $\gamma_{n}=\displaystyle\max_{x\in[0,1]}(a_{n}(x)-x)$. To this end, we recall the notation (5), i.e.,
@@
so that
@@
Therefore
@@
Now, we establish a Voronovskaja-type asymptotic formula for the operators $\mathscr{L}_{n}$.
@@
Since the function $\psi(y)$ is bounded and tends to zero as $y\to 0^{+}$, fixed $\varepsilon>0$, there exists $\tilde{\delta}>0$ such that, for every $y$ with $0<y\leq\tilde{\delta}$, then $|\psi(y)|<\varepsilon$. As a consequence, we can write
$n|R_{n}(x)|$ $\leq n\Bigg\{\sum_{\left|\frac{k}{n}-a_{n}(x)\right|\leq\frac{\tilde{\delta}}{2}}+\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\tilde{\delta}}{2}}\Bigg\}\left(\frac{k}{n}-x\right)^{2}\left|\psi\left(\frac{k}{n}-x\right)\right|\,$
@@
$\left|\frac{k}{n}-x\right|\leq\left|\frac{k}{n}-a_{n}(x)\right|+\left|a_{n}(x)-x\right|\leq\tilde{\delta}$,
@@
$R_{2}$ $\leq\|\psi\|_{\infty}\cdot n\ \sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\tilde{\delta}}{2}}\left(\frac{k}{n}-x\right)^{2}\,p_{n,k}(a_{n}(x))$
$\leq\|\psi\|_{\infty}\cdot n\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\tilde{\delta}}{2}}\left[2\left(\frac{k}{n}-a_{n}(x)\right)^{2}+2\left(a_{n}(x)-x\right)^{2}\right]\,p_{n,k}(a_{n}(x))$
$\leq{2}\|\psi\|_{\infty}\left\{n\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\tilde{\delta}}{2}}\left(\frac{k}{n}-a_{n}(x)\right)^{2}\,p_{n,k}(a_{n}(x))+n(a_{n}(x)-x)^{2}\right\},$
@@
$\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\tilde{\delta}}{2}}p_{n,k}(a_{n}(x))\leq C\,n^{-2}.$
@@
${2}\|\psi\|_{\infty}\,n\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\tilde{\delta}}{2}}\left(\frac{k}{n}-a_{n}(x)\right)^{2}p_{n,k}(a_{n}(x))$
$\leq{2}\|\psi\|_{\infty}\,n\sum_{\left|\frac{k}{n}-a_{n}(x)\right|>\frac{\tilde{\delta}}{2}}p_{n,k}(a_{n}(x))\leq{2}\|\psi\|_{\infty}C\,n^{-1}.$
@@ 5 Saturation results and inverse theorems
where $f\in C^{2}([0,1])$. By simple calculations, it is easy to see that the above differential operator can be rewritten in the following standard form:
@@
In [21] Garrancho and Cárdenas-Morales, inspired by the so-called parabola technique introduced by Bajsanski and Bojanic in [13] (see also Chapter 5 of [18] for a more detailed explanation), established the following useful result involving differential operators of the form (24).
@@ 6 Shape-preserving properties
In this section we study some shape preserving properties of the linear and positive operators $\mathscr{L}_{n}$.
@@
$\left(\frac{\mathscr{L}_{n}f}{\ln_{\mu}}\right)^{\prime\prime}(x)=$ $\hskip 2.84526pta_{n}^{\prime\prime}(x)\sum_{k=0}^{n-1}\Delta_{1}f_{\mu}\left(\frac{k}{n}\right)\,\,p_{n-1,k}(a_{n}(x))$
$+\frac{n-1}{n}a_{n}^{\prime}(x)^{2}\hskip 2.84526pt\sum_{k=0}^{n-2}\Delta_{2}f_{\mu}\left(\frac{k}{n}\right)\,\,p_{n-2,k}(a_{n}(x)).$
Since $a_{n}(x)$ is increasing on the interval $[0,1]$, it follows from expressions (27) and (28) that, if $f_{\mu}(x)$ is increasing on $[0,1]$, both ${\mathscr{L}_{n}f}/{\ln_{\mu}}$ and $\mathscr{L}_{n}f$ are increasing on $[0,1]$. And if in addition $f_{\mu}(x)$ is concave then, taking into account that $a_{n}(x)$ is concave, ${\mathscr{L}_{n}f}/{\ln_{\mu}}$ is concave as well.
@@
and $\mathcal{P}$ denotes the family of all the partitions of the interval $[0,1]$, that is, sets of the form $P=\{0=x_{0}<x_{1}<\ldots<x_{m}=1\}$. Furthermore, we recall that $BV_{\mu}([0,1])$ can be endowed with the norm
@@
Now, since the expression (27) holds, we obtain that
@@
(a) If $f_{\mu}(x)$ is increasing and convex, then, for every $n\in\mathbb{N}$ and $x\in]0,1[$,
@@
(b) If $f_{\mu}(x)$ is decreasing and concave, then, for every $n\in\mathbb{N}$ and $x\in]0,1[$,
@@
We will prove part $(a)$ since part $(b)$ follows immediately applying the same considerations of $(a)$ to the function f(x)-f(x), so that $f_{\mu}(x)$ will be increasing and convex.
@@
$\mathscr{L}_{n}(f,x)=\ln_{\mu}(x)\cdot B_{n}\left(f_{\mu},a_{n}(x)\right).$
Since $f_{\mu}(x)$ is convex on $[0,1]$, as a consequence of the classical shape preserving properties of Bernstein operators (see [19]), we have that
@@
for every $y\in[0,1]$. Moreover, since $a_{n}(x)\geq a_{n+1}(x)$ for every $x\in[0,1]$, $n\in\mathbb{N}$, and $f_{\mu}(x)$ is increasing, it follows that
@@
Finally, $\mathscr{L}_{n}(f,x)\to f(x)$, as $n\to+\infty$, on every continuity point $x$ of $f(x)$ (Theorem 3.1), therefore on every $x\in]0,1[$: indeed, $f_{\mu}(x)$ is continuous in ]0,1[]0,1[, since it is convex, and so is $f(x)=f_{\mu}(x)\cdot\ln_{\mu}(x)$. Therefore, since $\mathscr{L}_{n}(f,x)$ is decreasing with respect to $n$, we deduce that
@@ 7 Final remarks and future developments: a simple application to signal denoising
In this paper, we introduced a new definition of positive linear operators, which is characterized by the preservation of the function $\ln_{\mu}(x)$. For such operators, we established a number of approximation properties, such as uniform convergence, Voronovskaja-type formula, saturation and inverse theorems based on the solution of a suitable differential problem, and shape preserving properties.
@@
In real world situations, it can occur that a given signal $f$ is characterized by the presence of multiplicative noise, such as Gaussian multiplicative noise, uniform multiplicative noise, and several others. For instance, an example of multiplicative noise is the so-called speckle (e.g., [12, 37]), that typically affects remote sensing data, or tomography and ultrasound echography images. Other cases in which multiplicative noise arises is in stochastic models for finance or in wireless signal applications.
@@
Let $f:[0,1]\to\mathbb{R}^{+}$ be a given signal modelling a physical quantity. Suppose that $f$ is sampled (by a certain measurement tool) at a certain time $t>0$ in the uniformly spaced nodes $k/n$, k=0,1,,nk=0,1,...,n, $n\in\mathbb{N}$, obtaining a sequence of measured sample values $(y_{k,n})_{k}$.
We now suppose that $f$ is affected by a multiplicative noise function of the form: $n(x,t)=1+\mu(t)+x$, where the term $\mu(t)$ will be now viewed as a Gaussian random variable of mean 0 and variance $\sigma^{2}$, namely
@@
Hence, we expect that the sampled values $(y_{k,n})_{k}$, measured at time $t>0$, are of the form:
$y_{k,n}\approx\left(1+\mu(t)+\frac{k}{n}\right)\,f\left(\frac{k}{n}\right).$
Let now consider $t>0$ for which $\mu(t)>0$ and define $g(x):=\left(1+\mu(t)+x\right)f(x)$; thus
$\ln(g(x))\ =\ \ln_{\mu(t)}(x)\,+\,\ln(f(x)).$
Applying now the logarithmic operators ${\mathscr{L}}_{n}$ with $\mu(t)$ in place of the previously used parameter $\mu$, we get, by (2):
@@
where in (29) we replaced ${\mathscr{L}}_{n}(\ln g,x)$ with its approximated version computed by means of the sampled values $y_{k,n}$, namely
${\mathscr{L}}_{n}(\ln g,x)\ \approx\ \ln_{\mu(t)}(x)\sum_{k=0}^{n}\frac{\ln(y_{k,n})}{\ln_{\mu(t)}\left(\frac{k}{n}\right)}\hskip 2.84526ptp_{n,k}(a_{n}(x)).$
In conclusion, the final denoising formula is:
@@
We now present a numerical example. Let us consider, for instance, three values of $\mu(t)$ randomly generated as a Gaussian random variable ${\cal N}(0,\sigma^{2})$, with $\sigma=0.5$
$\mu_{1}:=\mu(t_{1})=0.2688,\quad\mu_{2}:=\mu(t_{2})=0.9169,\quad\mu_{3}:=\mu(t_{3})=1.1294,$
and let $f:[0,1]\to\mathbb{R}^{+}$ defined as
[…]
We now consider the following noise-free reconstructions of $f$, for $n=10,30$ in Fig.s 1, 2 and 3, implemented by MATLAB software. In the captions it is also reported the maximum reconstruction error computed between the original signal $f$ and the reconstructed one by means of (30) on the considered nodes.
Figure 1: The (blue) solid line represents the original signal $f$, the dotted (red) plot represents the signal $g$ affected by the multiplicative noise, the asterisk (yellow) plot represent the denoised signal by (30) for $\mu_{1}$, with $n=10$ (on the top) and $n=30$ (on the bottom). The maximum reconstruction errors are $0.1109$ and $0.0343$, respectively.
Figure 2: The (blue) solid line represents the original signal $f$, the dotted (red) plot represents the signal $g$ affected by the multiplicative noise, the asterisk (yellow) plot represent the denoised signal by (30) for $\mu_{2}$, with $n=10$ (on the top) and $n=30$ (on the bottom). The maximum reconstruction error are $0.0658$ and $0.0202$, respectively.
Figure 3: The (blue) solid line represents the original signal $f$, the dotted (red) plot represents the signal $g$ affected by the multiplicative noise, the asterisk (yellow) plot represent the denoised signal by (30) for $\mu_{3}$, with $n=10$ (on the top) and $n=30$ (on the bottom). The maximum reconstruction errors are $0.0622$ and $0.0191$, respectively.
Obviously, the above experiment represents only a toy model which serves to introduce a possible application of the new logarithmic-type operators; the main idea was to exploit the logarithm-preservation property besides the approximation properties of ${\mathscr{L}}_{n}$ to linearize a nonlinear source of noise. In future works, we aim to improve the above simple model, for instance, in order to deal with multivariate signals to implement a despeckle algorithm for remote sensing images (see, e.g. [11, 35, 39]).